OpenStudy (anonymous):

Find the slope of the line passing through the points (1/2,2/3) and (5/4, - 5/6) .

OpenStudy (anonymous):

\[\frac{y_1-y_2}{x_1-x_2}=\frac{\frac{-5}{6}-\frac{2}{3}}{\frac{5}{4}-\frac{1}{2}}\]

OpenStudy (anonymous):

Now simplify and you will have your answer :)

OpenStudy (anonymous):

-9/6 over 3/4??

OpenStudy (anonymous):

ok \[\frac{-9}{6} \div \frac{3}{4}=\frac{-9}{6} \cdot \frac{4}{3}\]

OpenStudy (anonymous):

Find that product there and boom! you will have it! You can simplify before multiplying

OpenStudy (anonymous):

The answer is -2?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

you deserve a medal! good job!

OpenStudy (anonymous):

Thanks =)